# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

### v2

@@ 8.3 $G$-linear semiorthogonal decompositions
Follows from [38, Corollary 2.3]. ∎
The goal of the rest of this section—and the two sections after it—is to construct a collection of objects $X_{p}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$, $p\in W$, each satisfying the condition in (2) of Proposition 8.8; by part (1) of the Proposition each of those objects will also be $B$-exceptional. Furthermore, the collection of objects $X_{p}$ will form a $G$-linear semiorthogonal decomposition of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$; the ultimate statement is Theorem 11.5.
8.5 Generating $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$
@@
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. The smallest strictly full triangulated subcategory of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ that contains the set of modules $\nabla_{\lambda}$, $\lambda\in X(T)_{+}$, is the whole $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$.
@@
First we show that ${\mathrm{rep}}(G)$ lies in the subcategory. We choose a real valued additive injective height function ${\mathrm{ht}}$ on the weight lattice which is positive on positive roots. Consider a nonzero $M\in{\mathrm{rep}}(G)$. Say $\mu$ is its highest weight with respect to ${\mathrm{ht}}$. Thus ${\mathrm{ht}}(\mu)\geq 0$. Assume that all representations with a smaller highest weight are in the subcategory. We have a map $\Delta_{\mu}\to\nabla_{\mu}$ whose kernel and cokernel have lower weights, so $\Delta_{\mu}$ lies in the subcategory. Let $M_{\mu,\mathrm{triv}}$ denote the weight space $M_{\mu}$ provided with a trivial $G$ action. By [23, Proposition 21] there is a natural map $\Delta_{\mu}\otimes_{\sf k}M_{\mu,\mathrm{triv}}\to M$ whose kernel and cokernel have lower weights. Now notice that $\Delta_{\mu}\otimes_{\sf k}M_{\mu,\mathrm{triv}}$ lies in the subcategory. So $M$ does too. So ${\mathrm{rep}}(G)$ lies in the subcategory. Now use that every object of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ is quasi-isomorphic to a bounded complex in ${\mathrm{rep}}(G)$ and that bounded complexes are repeated cones of pure objects. (cf. “Stupid truncations” [33, 2.5].) ∎
@@
Let $S$ be a set of objects of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$.
Then ${\mathrm{hull}}(\{\nabla_{\nu}\otimes M\mid\nu\in X(T)_{+},M\in S\})$ is $G$-linear.
@@
[…]
8.6 Generating $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$
@@
Given a $p\in W$, the triangulated hull in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{Rep}}(B))$ of the two categories
${\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succ p,\lambda\in X(T)_{+}}),$
${\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(T)_{+}})$
is $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$.
@@
This will take the rest of the section. ∎
8.7 Generating ${\mathrm{rep}}(B)$
@@ 8.8 The set $\{M_{v}\}_{v\in W}$
Let us be given a set $\{M_{v}\}_{v\in W}$ of objects of ${\mathrm{rep}}_{\mathrm{fr}}(B)$ with the following properties. The multiplicity of the weight $e_{v}$ in $M_{v}$ is one. Every weight $\lambda$ of $M_{v}$ satisfies $(\lambda,\lambda)\leq(e_{v},e_{v})$. If $\lambda$ is a weight of $M_{v}$ with $(\lambda,\lambda)=(e_{v},e_{v})$, then $\lambda$ is a weight of $P(-e_{v})^{*}$. So all weights $\lambda$ of $M_{v}$ precede $e_{v}$ in the antipodal excellent order, notation $\lambda\leq_{a}e_{v}$. Examples of possible choices of $M_{v}$ are ${\sf k}_{e_{v}}$, $Q(e_{v})$, $P(-e_{v})^{*}$. We use $e^{\lambda}$ as another notation for ${\sf k}_{\lambda}$.
@@
is the category ${\mathrm{rep}}(B)$ of finite dimensional $B$-modules.
We first prove two propositions.
@@
Put $\tau=\varpi-\omega_{j}$. Let $W_{I}$, $W_{J}$, $W_{K}$ be the stabilizers in $W$ of $\tau$, $\omega_{j}$, $\varpi$ respectively. Choose $z$ minimal so that $w^{-1}\mu=\tau+z\omega_{j}$. Assume $(\mu,\mu)\geq(\varpi,\varpi)$. Choose a reduced expression $s_{k}s_{k-1}\cdots s_{1}$ for $z$ and put $z_{0}={\mathrm{id}}$, $z_{i}=s_{i}z_{i-1}$. Observe for $0<i\leq k$ that $z_{i-1}\omega_{j}$ is a weight of $P(z_{i}\omega_{j})$, which has socle of weight $z_{i}\omega_{j}$. The socle must lie in the $B$-submodule generated by the weight space of weight $z_{i-1}\omega_{j}$. This implies $z_{i}\omega_{j}\leq_{d}z_{i-1}\omega_{j}$, where $\leq_{d}$ is the dominance partial order, defined in 2.1. The path $z_{0}\omega_{j},\cdots,z_{k}\omega_{j}$ from $\omega_{j}$ to $z\omega_{j}$ is thus strictly descending for $\leq_{d}$. Along the path $(\tau,z_{i}\omega_{j})$ can only go down, but $(\tau,u\omega_{j})<(\tau,v\omega_{j})$ implies $(\tau+u\omega_{j},\tau+u\omega_{j})<(\tau+v\omega_{j},\tau+v\omega_{j})$. So the path must consist of steps in directions perpendicular to $\tau$. So $z\in W_{I}$. Now $w$ is minimal in its coset $wW_{K}$, and moreover we have $w\leq ws_{j}$. That makes it minimal in $wW_{I}$ also. This shows $wz\geq w$. Now note that $\mu=w(\tau+z\omega_{j})=w(z\tau+z\omega_{j})=wz\varpi$. ∎
@@
We claim that then $e^{\lambda}$ is in the subcategory. The theorem easily follows from the claim by induction along a suitable order on weights.
@@
The second case is that some $\alpha_{j}^{\vee}(\varpi)>1$. Put $\tau=\varpi-\omega_{j}$. Again $\tau$ is strictly shorter than $\varpi$. By Proposition 8.18 the weights $\mu$ of $N=e^{w\tau}\otimes\nabla(\omega_{j})$ are either shorter than $\lambda$ or they are of the form $v\varpi$ with $v\geq w$. So $e^{\mu}$ is in the subcategory for all weights of $N$ different from $\lambda$ and $\lambda$ has multiplicity one in $N$. It follows that $e^{\lambda}$ is in the subcategory.
@@
Let ${\sf k}$ be an algebraically closed field. In the Steinberg basis $\{{\sf k}_{e_{v}}\}_{v\in W}$ of $R(B)$ over $R(G)$, one may replace some ${\sf k}_{e_{v}}$ by $Q(e_{v})$ and some other ${\sf k}_{e_{w}}$ by $P(-e_{w})^{*}$. The result will still be a basis.
@@
The results of this section give an alternative proof of Steinberg’s theorem [44] that the ${\sf k}_{e_{w}}$ generate $R(B)$ over $R(G)$. See also Corollary 10.2. Our proof assumes only that $G$ is split and ${\sf k}$ is a field or ${\mathbb{Z}}$, while Steinberg takes the field algebraically closed. But by [43, §3] this makes no difference.
@@
By Theorem 8.16, for any $p\in W$ the set $\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succ p,\lambda\in X(T)_{+}}\cup\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(T)_{+}}$ generates ${\mathrm{rep}}(B)$ as an abelian category. Considering the canonical truncation of an object of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ and resolving each cohomology via the above generators of ${\mathrm{rep}}(B)$, we get the statement. One could also use stupid truncations, cf. [33, 2.5]. ∎
@@ 9 Construction of the objects $X_{p}$ and $Y_{p}$
${\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succ p,\lambda\in X(T)_{+}}\cup\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(T)_{+}})=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B)).$ (18)
@@ 9.1 Cut at $p\in W$
  • ${\sf Q}_{\succeq p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succeq p,\lambda\in X(T)_{+}})$
  • ${\sf Q}_{\succ p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succ p,\lambda\in X(T)_{+}})$
  • ${\sf P}_{\preceq p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(T)_{+}})$
  • ${\sf P}_{\prec p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\prec p,\lambda\in X(T)_{+}})$
Denote $\mathcal{D}:={\mathrm{hull}}({\sf Q}_{\succ p}\cup{\sf P}_{\preceq p})$. Then (18) gives that $\mathcal{D}$ is just $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$.
@@
By Corollary 6.7, for all $p\in W$:
${\sf Q}_{\succ p}\subseteq{\sf P}_{\preceq p}^{\perp}.$ (20)
@@
Thus, by Definition 14, $\mathcal{D}$ has a semiorthogonal decomposition:

### v3

@@ 8.3 $\mathbf{G}$-linear semiorthogonal decompositions
Follows from [41, Corollary 2.3]. ∎
The goal of the rest of this section—and the two sections after it—is to construct a collection of objects $X_{p}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$, $p\in W$, each satisfying the condition in (2) of Proposition 8.11; by part (1) of the Proposition each of those objects will also be $\mathbf{B}$-exceptional. Furthermore, the collection of objects $X_{p}$ will form a $\mathbf{G}$-linear semiorthogonal decomposition of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))$; the ultimate statement is Theorem 11.6.
8.5 Generating $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$
@@
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. The smallest strictly full triangulated subcategory of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$ that contains the set of modules $\nabla_{\lambda}$, $\lambda\in X(\mathbf{T})_{+}$, is the whole $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$.
@@
First we show that ${\mathrm{rep}}(\mathbf{G})$ lies in the subcategory. We choose a real valued additive injective height function ${\mathrm{ht}}$ on the weight lattice which is positive on positive roots. Consider a nonzero $M\in{\mathrm{rep}}(\mathbf{G})$. Say $\mu$ is its highest weight with respect to ${\mathrm{ht}}$. Thus ${\mathrm{ht}}(\mu)\geq 0$. Assume that all representations with a smaller highest weight are in the subcategory. We have a map $\Delta_{\mu}\to\nabla_{\mu}$ whose kernel and cokernel have lower weights, so $\Delta_{\mu}$ lies in the subcategory. Let $M_{\mu,\mathrm{triv}}$ denote the weight space $M_{\mu}$ provided with a trivial $\mathbf{G}$ action. By [26, Proposition 21] there is a natural map $\Delta_{\mu}\otimes_{\sf k}M_{\mu,\mathrm{triv}}\to M$ whose kernel and cokernel have lower weights. Now notice that $\Delta_{\mu}\otimes_{\sf k}M_{\mu,\mathrm{triv}}$ lies in the subcategory. So $M$ does too. So ${\mathrm{rep}}(\mathbf{G})$ lies in the subcategory. Now use that every object of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$ is quasi-isomorphic to a bounded complex in ${\mathrm{rep}}(\mathbf{G})$ and that bounded complexes are repeated cones of pure objects. (cf. “Stupid truncations” [36, 2.5].) ∎
@@
Let $S$ be a set of objects of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$.
Then ${\mathrm{hull}}(\{\nabla_{\nu}\otimes M\mid\nu\in X(\mathbf{T})_{+},M\in S\})$ is $\mathbf{G}$-linear.
@@
By Proposition 8.17 we have ${\mathrm{hull}}(\{\nabla_{\nu}\mid\nu\in X(\mathbf{T})_{+}\})=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$. So ${\mathrm{hull}}(\{\nabla_{\nu}\otimes M\mid\nu\in X(\mathbf{T})_{+},M\in S\})$ equals ${\mathrm{hull}}(\{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))\otimes M\mid M\in S\})$. ∎
8.6 Generating $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))$
@@
Given a $p\in W$, the triangulated hull in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{Rep}}(\mathbf{B}))$ of the two categories
${\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succ p,\lambda\in X(\mathbf{T})_{+}}),$
${\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(\mathbf{T})_{+}})$
is $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))$.
@@
This will take the rest of the section. We will categorify the theorem of Steinberg that says that the $[{\sf k}_{e_{v}}]$ generate $R(\mathbf{B})$ as an $R(\mathbf{G})$-module and we will apply the same reasoning with a few ${\sf k}_{e_{v}}$’s replaced with $Q(e_{v})$ or $P(-e_{v})^{*}$. Our arguments are similar to the proof of [3, Theorem 2]. ∎
8.7 Generating ${\mathrm{rep}}({\mathbf{B}})$
@@ 8.8 The set $\{M_{v}\}_{v\in W}$
Let us be given a set $\{M_{v}\}_{v\in W}$ of objects of ${\mathrm{rep}}_{\mathrm{fr}}({\mathbf{B}})$ with the following properties. The multiplicity of the weight $e_{v}$ in $M_{v}$ is one. Every weight $\lambda$ of $M_{v}$ satisfies $(\lambda,\lambda)\leq(e_{v},e_{v})$. If $\lambda$ is a weight of $M_{v}$ with $(\lambda,\lambda)=(e_{v},e_{v})$, then $\lambda$ is a weight of $P(-e_{v})^{*}$. So all weights $\lambda$ of $M_{v}$ precede $e_{v}$ in the antipodal excellent order, notation $\lambda\leq_{a}e_{v}$. Examples of possible choices of $M_{v}$ are ${\sf k}_{e_{v}}$, $Q(e_{v})$, $P(-e_{v})^{*}$. We use $e^{\lambda}$ as another notation for ${\sf k}_{\lambda}$.
@@
is the category ${\mathrm{rep}}({\mathbf{B}})$ of finite dimensional ${\mathbf{B}}$-modules.
We first prove a lemma and two propositions.
Lemma 8.21.
Let $\lambda\in X(\mathbf{T})$, $\alpha\in\Pi$. Then $s_{\alpha}\lambda\leq_{e}\lambda$ if and only if $s_{\alpha}\lambda\geq_{d}\lambda$.
Recall that $s_{\alpha}\lambda\leq_{e}\lambda$ if and only if $s_{\alpha}\lambda\geq_{a}\lambda$ by Lemma 3.10.
Proof.
Write $\lambda=w\nu$ with $\nu$ dominant and $w$ minimal. We have $s_{\alpha}w>w$ or $s_{\alpha}w<w$. If $s_{\alpha}w>w$, then $\lambda$ is a weight of $P(s_{\alpha}\lambda)$ and the socle of $P(s_{\alpha}\lambda)$, which has weight $s_{\alpha}\lambda$ by [49, Lemma 2.2.9], must lie in the $\mathbf{B}$-submodule of $P(s_{\alpha}\lambda)$ generated by the weight space of weight $\lambda$. This implies $s_{\alpha}\lambda\leq_{d}\lambda$. The other case is similar. ∎
@@
Put $\tau=\varpi-\omega_{j}$. Let $W_{I}$, $W_{J}$, $W_{K}$ be the stabilizers in $W$ of $\tau$, $\omega_{j}$, $\varpi$ respectively. Choose $z$ minimal so that $w^{-1}\mu=\tau+z\omega_{j}$. Assume $(\mu,\mu)\geq(\varpi,\varpi)$. Choose a reduced expression $s_{k}s_{k-1}\cdots s_{1}$ for $z$ and put $z_{0}={\mathrm{id}}$, $z_{i}=s_{i}z_{i-1}$. As $z_{i+1}\omega_{j}>_{e}z_{i}\omega_{j}$, we have by Lemma 8.21 that the path $z_{0}\omega_{j},\cdots,z_{k}\omega_{j}$ from $\omega_{j}$ to $z\omega_{j}$ is strictly descending for $\leq_{d}$. Along the path $(\tau,z_{i}\omega_{j})$ can only go down, but $(\tau,u\omega_{j})<(\tau,v\omega_{j})$ implies $(\tau+u\omega_{j},\tau+u\omega_{j})<(\tau+v\omega_{j},\tau+v\omega_{j})$. So the path must consist of steps in directions perpendicular to $\tau$. So $z\in W_{I}$. Now $w$ is minimal in its coset $wW_{K}$, and moreover we have $w\leq ws_{j}$. That makes it minimal in $wW_{I}$ also. This shows $wz\geq w$. Now note that $\mu=w(\tau+z\omega_{j})=w(z\tau+z\omega_{j})=wz\varpi$. ∎
@@
We claim that then $e^{\lambda}$ is in the subcategory. The theorem easily follows from the claim by induction along $<_{a}$.
@@
The second case is that some $\alpha_{j}^{\vee}(\varpi)>1$. Put $\tau=\varpi-\omega_{j}$. Again $\tau$ is strictly shorter than $\varpi$. By Proposition 8.23 the weights $\mu$ of $N=e^{w\tau}\otimes\nabla(\omega_{j})$ are either shorter than $\lambda$ or they are of the form $v\varpi$ with $v\geq w$. So $e^{\mu}$ is in the subcategory for all weights of $N$ different from $\lambda$ and $\lambda$ has multiplicity one in $N$. It follows that $e^{\lambda}$ is in the subcategory. ∎
@@
If $\sigma$ is not a Steinberg weight, then there is a fundamental weight $\omega_{j}$ and a weight $\mu$ with $(\mu,\mu)<(\sigma,\sigma)$, so that the $\mathbf{B}$-module $L=\nabla_{\omega_{j}}\otimes e^{\mu}$ has $\sigma$ as a weight of multiplicity one, and so that all weights $\nu$ of $L$ satisfy $\nu\leq_{a}\sigma$.
Proof.
See the proof of Theorem 8.20. ∎
Corollary 8.25.
Let ${\sf k}$ be an algebraically closed field. Let $W$ be partitioned arbitrarily into three subsets $W_{1}$, $W_{2}$, $W_{3}$. One gets a basis of $R({\mathbf{B}})$ over $R(\mathbf{G})$ by taking $\{{\sf k}_{e_{v}}\}$ for ${v\in W_{1}}$, $Q(e_{v})$ for ${v\in W_{2}}$, $P(-e_{v})^{*}$ for ${v\in W_{3}}$.
@@
The results of this section give an alternative proof of Steinberg’s theorem [46] that the ${\sf k}_{e_{w}}$ generate $R({\mathbf{B}})$ over $R(\mathbf{G})$. See also Corollary 10.2 or [3, Theorem 2]. Our proof assumes only that $\mathbf{G}$ is split and ${\sf k}$ is a field or ${\mathbb{Z}}$, while Steinberg takes the field algebraically closed. But by [45, §3] this makes no difference.
@@
By Theorem 8.20, for any $p\in W$ the set $\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succ p,\lambda\in X(\mathbf{T})_{+}}\cup\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(\mathbf{T})_{+}}$ generates ${\mathrm{rep}}({\mathbf{B}})$ as an abelian category in a specific manner, corresponding in the derived category with taking cones, shifts, or using the $\mathbf{G}$-linear structure of Proposition 8.18. Therefore the hull contains all pure objects (objects concentrated in one degree). Considering the canonical truncation of an object of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))$ and resolving each cohomology via pure objects, we get the statement. One could also use stupid truncations, cf. [36, 2.5]. ∎
@@ 9 Construction of the objects $X_{p}$ and $Y_{p}$
${\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succ p,\lambda\in X(\mathbf{T})_{+}}\cup\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(\mathbf{T})_{+}})=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B})).$ (21)
@@ 9.1 Cut at $p\in W$
  • ${\sf Q}_{\succeq p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succeq p,\lambda\in X(\mathbf{T})_{+}})$
  • ${\sf Q}_{\succ p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{v\succ p,\lambda\in X(\mathbf{T})_{+}})$
  • ${\sf P}_{\preceq p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,\lambda\in X(\mathbf{T})_{+}})$
  • ${\sf P}_{\prec p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\prec p,\lambda\in X(\mathbf{T})_{+}})$
Remark 9.1.
We will need to find exceptional objects to conclude that ${\sf Q}_{\succeq p}$, ${\sf P}_{\preceq p}$, ${\sf Q}_{\succeq p}\cap{\sf P}_{\preceq p}$ are admissible subcategories. That is what this section and the next one are about.
Remark 9.2.
Do not confuse ${\sf Q}_{\succ p}$ with ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes Q(e_{v})\}_{e_{v}>_{a}e_{p},\lambda\in X(\mathbf{T})_{+}})$. The latter may seem more natural, as the $Q(\mu)$ belong with $>_{a}$, not $\succ$. The Steinberg weights show two faces, and this is essential. On the one hand they are indexed by the Weyl group and are thus ordered by the Bruhat order, on the other hand they can be characterised in terms of $>_{a}$, see Remark 10.3.
Denote $\mathcal{D}:={\mathrm{hull}}({\sf Q}_{\succ p}\cup{\sf P}_{\preceq p})$. Then (21) gives that $\mathcal{D}$ is just $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$.
@@
By Corollary 6.7 and Corollary 8.3, for all $p\in W$:
${\sf Q}_{\succ p}\subset{\sf P}_{\preceq p}^{\perp}.$ (23)
@@
Thus, by Definition 7.7, $\mathcal{D}$ has a semiorthogonal decomposition: